# Effect of Off-diagonal NSI Parameters on Entanglement Measurements in Neutrino Oscillations

### v1

@@ Abstract
In this work, we explore the influence of off-diagonal non-standard interaction (NSI) parameters on quantum entanglement within the three-flavor neutrino oscillation framework. By expressing three key entanglement measures: Entanglement of Formation (EOF), Concurrence, and Negativity in terms of oscillation probabilities, we analyze how these quantum correlations are affected by the NSI parameters $\epsilon_{e\mu}$, $\epsilon_{e\tau}$, and $\epsilon_{\mu\tau}$, including their complex phases. Using the DUNE experiment as a benchmark, we find that NSI effects are most significant at low energies, with Negativity showing the highest sensitivity even at high energies. It is observed that $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ affect entanglement measures mainly through the appearance channel, while the impact of $\epsilon_{\mu\tau}$ on EOF, Concurrence, and Negativity is predominantly linked to the disappearance channel.
@@ I Introduction
Neutrino oscillation Fukuda et al. [Super-Kamiokande Collaboration] (1998); Ahmad et al. [SNO Collaboration] (2002) is a quantum mechanical phenomenon that arises from the mixing between neutrino flavor and mass eigenstates. Since neutrino flavor states are quantum superpositions of mass eigenstates, a neutrino originated with a specific flavor can be detected as a different flavor during propagation. The probabilities of such oscillations are governed by six key parameters: three mixing angles $(\theta_{12},\theta_{13},\theta_{23})$, two mass-squared differences $(\Delta m^{2}_{21},\Delta m^{2}_{31})$, and a CP-violating phase, $\delta_{CP}$. Over the past few decades, a broad range of solar, atmospheric, reactor, and accelerator neutrino experiments have measured most of these parameters with remarkable precision. Despite this progress, several fundamental questions remain unresolved, most notably are the ordering of neutrino masses, commonly known as mass hierarchy problem, the exact value of the CP-violating phase $\delta_{CP}$, and whether the atmospheric mixing angle $\theta_{23}$ lies in the lower or upper octant. Addressing these unknowns is a primary goal of upcoming and ongoing high-precision neutrino oscillation experiments Abe et al. [T2K Collaboration] (2023); Acero et al. [NOvA Collaboration] (2022); Abi et al. [DUNE Collaboration] (2020); collaboration et al. (2022), which are designed not only to rigorously test the three-flavor paradigm but also to search for signatures of physics Beyond the Standard Model (BSM) Abi et al. (2021); Majhi et al. (2023); Mohanta (2025); Panda and Mohanta (2024); Panda et al. (2025); Pusty et al. (2024); Bera et al. (2025); Singha et al. (2024).
@@
Neutrino Physics has now entered the precision era, and at this point, it is crucial to study the sub-leading effects in neutrino oscillations, commonly termed as Non-Standard neutrino Interactions (NSIs), which introduce new coupling parameters to the theory, and thus play a pivotal role on the determination of unknowns in neutrino sector. A model-independent framework for exploring NSI effects in neutrino oscillations is discussed in Refs.Ohlsson (2013); Miranda and Nunokawa (2015); Farzan and Tortola (2018). Among the various possible new physics effects in the neutrino sector, NSIs have received significant attention and effectively modify neutrino production, detection, and most significantly propagation of neutrinos through matter Biggio et al. (2009). These interactions can be described by dimension-six four-fermion operators, which introduce additional terms into the neutrino evolution Hamiltonian. In this work, we focus on propagation NSI, as constraints on production and detection NSI are typically more stringent. The matter NSI terms are parametrized by dimensionless coefficients $\epsilon_{\alpha\beta}$ ($\alpha,\beta=e,\mu,\tau$) that modify the effective matter potential experienced by neutrinos in propagation. These include both diagonal (flavor-conserving) and off-diagonal (flavor-violating) components, where the off-diagonal terms can be complex and introduce new CP-violating effects. Since NSIs affect the coherent forward scattering of neutrinos with electrons and quarks in matter, they can lead to significant modifications in oscillation probabilities, especially in long-baseline experiments with specific beamlines and detector setups Deepthi et al. (2015); Fukasawa et al. (2017); Masud et al. (2016); de Gouvêa and Kelly (2016); Deepthi et al. (2018); Blennow et al. (2017); Agarwalla et al. (2016); Masud and Mehta (2016); Coloma and Schwetz (2017); Liao et al. (2016); Coloma (2016).
@@
Recent developments in quantum information theory have opened a novel and exciting perspective in the neutrino system Alok et al. (2016); Banerjee et al. (2015); Formaggio et al. (2016); Fu and Chen (2017); Naikoo et al. (2019, 2020); Shafaq and Mehta (2021); Sarkar and Dixit (2021); Blasone et al. (2023); Chattopadhyay and Dighe (2023); Konwar et al. (2024); Dixit et al. (2024); Bouri et al. (2024); Banerjee et al. (2024); Konwar and Yadav (2025); Konwar (2025). Specifically, the concept of quantum entanglement, once considered relevant primarily in systems like atomic and optical systems, is now being investigated in the context of fundamental particles, including neutrinos Blasone et al. (2009, 2010, 2021); Kumar Jha et al. (2021); Konwar and Yadav (2024); Banerjee et al. (2024). Entanglement represents a core feature of quantum mechanics, characterising non-classical correlations between different components of a system Horodecki et al. (2009). In neutrino physics, the coherent superposition of mass and flavor eigenstates naturally leads to a form of single-particle mode entanglement Blasone et al. (2008, 2009), where the different flavor modes of a single neutrino become quantum-mechanically correlated. The key insight is that these entanglement measures can be quantified using neutrino oscillation probabilities, which are directly measurable in neutrino oscillation experiments. To quantify entanglement in neutrino systems, several well-established measures from quantum information theory are employed. In this work, we focus on three such measures: englement of Formation (EOF) Bennett et al. (1996); Guo and Zhang (2020), Concurrence Hill and Wootters (1997); Wootters (1998); Guo and Gour (2019) and Negativity Peres (1996); Vidal and Werner (2002); Sabín and García-Alcaine (2008).
@@
Among the various entanglement measures, the EOF plays a prominent role and captures the minimum average entanglement over all pure state decompositions of a mixed state. Concurrence, initially formulated for bipartite qubit systems, has been successfully extended to tripartite scenarios and provides a robust way to evaluate the degree of entanglement through the reduced density matrices. Another widely used measure is negativity, which is derived from the Peres-Horodecki criterion Peres (1996). It quantifies entanglement based on the partial transpose of the density matrix, acquiring negative eigenvalues, which serves as a clear signature of non-separability and, consequently, the presence of entanglement in the system. All three measures can be effectively expressed in terms of flavor transition probabilities, allowing the entanglement to be quantified using experimentally observable quantities and providing a framework to study quantum correlations in neutrino oscillations. In this study, we explore the behavior of three key entanglement measures within the framework of three-flavor neutrino oscillations by formulating them in terms of oscillation probabilities. We further examine how these measures are influenced by the presence of three complex off diagonal NSI scenarios.
@@
The paper is organized as follows. Section II presents the theoretical formalism underlying neutrino oscillation with NSI, along with a description of mode entanglement in the three-flavor neutrino system. In Section III, we define the three entanglement measures used in our analysis: EOF, Concurrence and Negativity, and express them in terms of neutrino oscillation probabilities. Section IV discusses the simulation setup based on the DUNE experiment. Section V emphasis the major results coming out from our work. Finally, Section VI summarizes the conclusions of the study.
@@ II Formalism
Neutrino oscillations arise as a consequence of both neutrino mixing and non-degenerate neutrino masses. The neutrino mass eigenstates |$\ket{\nu_{j}},j=1,2,3$, associated with definite masses $m_{j}$ and energies $E_{j}$, evolve in time of the form:
|$\ket{\nu_{j}(t)}=e^{-\iota E_{j}t}\ket{\nu_{j}}.$ (1)
The flavor eigenstates |$\ket{\nu_{\alpha}}$, $\alpha=e,\mu,\tau$ are linear superpositions of the mass eigenstates, and their time evolution is governed by the unitary operator $U_{f}(t)$, such that
|$\ket{\nu_{\alpha}(t)}=U_{f}(t)\ket{\nu_{\beta}},$ (2)
where $U_{f}(t)=Ue^{-\iota H_{m}t}U^{-1}$ is the time-evolution operator with $\mathcal{H}_{m}=\text{diag}(E_{1},E_{2},E_{3})$ and $U$ representing the PMNS matrix that relates the flavor and mass eigenstates. The flavor state |$\ket{\nu_{\alpha}(t)}$ at time $t$ can then be expressed in terms of the flavor basis as:
|$\ket{\nu_{\alpha}(t)}={U}_{\alpha e}(t)\ket{\nu}_{e}+{U}_{\alpha\mu}(t)\ket{\nu}_{\mu}+{U}_{\alpha\tau}(t)\ket{\nu}_{\tau}.$ (3)
The transition probability for a neutrino of flavor $\alpha$ to be detected as a neutrino of flavor $\beta$ at time $t$ is given by:
$P_{(\nu_{\alpha}\rightarrow\nu_{\beta})}(t)=\left|\left<\nu_{\beta}|\nu_{\alpha}(t)\right>\right|^{2}=\left|U_{\alpha\beta}(t)\right|^{2}$ (4)
In the scenarios beyond the SM, neutrino propagation can be affected by NSI with matter, which introduce additional effective potentials modifying the flavor evolution Hamiltonian. The NSI formalism extends the standard oscillation framework by incorporating these new interactions as perturbations in the matter potential, typically parameterized by dimensionless coefficients $\epsilon_{\alpha\beta}$ with $\alpha,\beta=e,\mu,\tau$ Ohlsson (2013); Miranda and Nunokawa (2015); Liao et al. (2017). This approach allows for a systematic study of how beyond-Standard-Model effects can influence neutrino oscillation probabilities.
@@
Here, each term in this expression represents the occupation of one specific flavor mode (e, $\mu$, or $\tau$) while the others remain unoccupied.
To analyze this entanglement quantitatively, the density matrix of the system at time $t$ is constructed as, $\rho_{ABC}^{\alpha}(t)=\ket{\nu_{\alpha}(t)}\bra{\nu_{\alpha}(t)}$ |. For the neutrino state, the density matrix takes the form Banerjee et al. (2024):
${}\rho_{ABC}^{\alpha}(t)=\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&\rho_{22}^{\alpha}&\rho_{23}^{\alpha}&0&\rho_{25}^{\alpha}&0&0&0\\ 0&\rho_{32}^{\alpha}&\rho_{33}^{\alpha}&0&\rho_{35}^{\alpha}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&\rho_{52}^{\alpha}&\rho_{53}^{\alpha}&0&\rho_{55}^{\alpha}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{pmatrix},$ (6)
where the elements of this matrix can be represented as.
$\rho_{22}^{\alpha}=\left|\bar{U}_{\alpha\tau}(t)\right|^{2};~{}~{}~{}~{}~{}~{}~{}~{}\rho_{23}^{\alpha}=\bar{U}_{\alpha\tau}(t)\bar{U}_{\alpha\mu}^{\ast}(t);~{}~{}~{}~{}\rho_{25}^{\alpha}=\bar{U}_{\alpha\tau}(t)\bar{U}_{\alpha e}^{\ast}(t);$
$\rho_{32}^{\alpha}=\bar{U}_{\alpha\mu}(t)\bar{U}_{\alpha\tau}^{\ast}(t);~{}~{}~{}\rho_{33}^{\alpha}=\left|\bar{U}_{\alpha\mu}(t)\right|^{2};~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho_{35}^{\alpha}=\bar{U}_{\alpha\mu}(t)\bar{U}_{\alpha e}^{\ast}(t);$
$\rho_{52}^{\alpha}=\bar{U}_{\alpha e}(t)\bar{U}_{\alpha\tau}^{\ast}(t);~{}~{}~{}~{}\rho_{53}^{\alpha}=\bar{U}_{\alpha e}(t)\bar{U}_{\alpha\mu}^{\ast}(t);~{}~{}~{}~{}\rho_{55}^{\alpha}=\left|\bar{U}_{\alpha e}(t)\right|^{2}.$ (7)
@@
The probabilities are as follows: $P_{\alpha e}(t)=\left|\bar{U}_{\alpha e}(t)\right|^{2}$, $P_{\alpha\mu}(t)=\left|\bar{U}_{\alpha\mu}(t)\right|^{2}$, and $P_{\alpha\tau}(t)=\left|\bar{U}_{\alpha\tau}(t)\right|^{2}$. In the next section, we will discuss the mathematical expressions of EOF, Concurrence and Negativity in three-flavor neutrino system.

### v2

@@ Abstract
In this work, we explore the influence of off-diagonal non-standard interaction (NSI) parameters on quantum entanglement within the three-flavor neutrino oscillation framework. By expressing three key entanglement measures: Entanglement of Formation (EOF), Concurrence, and Negativity in terms of oscillation probabilities, we analyze how these quantum correlations are affected by the NSI parameters $\epsilon_{e\mu}$, $\epsilon_{e\tau}$, and $\epsilon_{\mu\tau}$, including their complex phases. The quantum correlation measures considered in this work cannot be extracted directly from event rates, but solely in terms of oscillation probabilities. Using the DUNE experiment as a reference point, our analysis shows that NSI effects are most pronounced at lower energies, while Negativity continuing to dominate even at higher energies. It is observed that $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ affect entanglement measures mainly through the appearance channel, while the impact of $\epsilon_{\mu\tau}$ on EOF, Concurrence, and Negativity is predominantly linked to the disappearance channel. Further, our results show that Negativity is more sensitive than EOF and Concurrence in the [Energy ($E$) – $\delta_{CP}$] plane under the influence of off-diagonal NSI scenarios, displaying a clear dependence of the CP-violating phase, $\delta_{CP}$ on specific energy ranges, particularly in the lower energy regime.
@@ I Introduction
Neutrino oscillation Fukuda et al. [Super-Kamiokande Collaboration] (1998); Ahmad et al. [SNO Collaboration] (2002) is a quantum mechanical phenomenon that arises from the mixing between neutrino flavor and mass eigenstates. Since neutrino flavor states are quantum superpositions of mass eigenstates, a neutrino originated with a specific flavor can be detected as a different flavor during propagation. The probabilities of such oscillations are governed by six key parameters: three mixing angles $(\theta_{12},\theta_{13},\theta_{23})$, two mass-squared differences $(\Delta m^{2}_{21},\Delta m^{2}_{31})$, and a CP-violating phase, $\delta_{CP}$. Over the past few decades, a broad range of solar, atmospheric, reactor, and accelerator neutrino experiments have measured most of these parameters with remarkable precision. Despite this progress, several fundamental questions remain unresolved, most notably the ordering of neutrino masses, commonly known as the mass hierarchy problem, the exact value of the CP-violating phase $\delta_{CP}$, and whether the atmospheric mixing angle $\theta_{23}$ lies in the lower or upper octant. Addressing these unknowns is a primary goal of upcoming and ongoing high-precision neutrino oscillation experiments Abe et al. [T2K Collaboration] (2023); Acero et al. [NOvA Collaboration] (2022); Abi et al. [DUNE Collaboration] (2020); Abusleme et al. [JUNO Collaboration] (2022), which are designed not only to rigorously test the three-flavor paradigm but also to search for signatures of physics Beyond the Standard Model (BSM) Abi et al. [DUNE Collaboration] (2021); Majhi et al. (2023); Mohanta (2025); Panda and Mohanta (2025); Panda et al. (2025); Pusty et al. (2024); Bera et al. (2025); Singha et al. (2024).
@@
Neutrino Physics has now entered the precision era, and at this point, it is crucial to study the sub-leading effects in neutrino oscillations, commonly termed as Non-Standard neutrino Interactions (NSIs), which introduce new coupling parameters to the theory, and thus play a pivotal role on the determination of unknowns in the neutrino sector. A model-independent framework for exploring NSI effects in neutrino oscillations is discussed in Refs.Ohlsson (2013); Miranda and Nunokawa (2015); Farzan and Tortola (2018). Among the various possible new physics effects in the neutrino sector, NSIs have received significant attention and effectively modify neutrino production, detection, and most significantly propagation of neutrinos through matter Biggio et al. (2009). These interactions can be described by dimension-six four-fermion operators, which introduce additional terms into the neutrino evolution Hamiltonian. In this work, we focus on propagation NSI, as constraints on production and detection NSI are typically more stringent. The matter NSI terms are parametrized by dimensionless coefficients $\epsilon_{\alpha\beta}$ ($\alpha,\beta=e,\mu,\tau$) that modify the effective matter potential experienced by neutrinos in propagation. These include both diagonal (flavor-conserving) and off-diagonal (flavor-violating) components, where the off-diagonal terms can be complex and introduce new CP-violating effects. Since NSIs affect the coherent forward scattering of neutrinos with electrons and quarks in matter, they can lead to significant modifications in oscillation probabilities, especially in long-baseline experiments with specific beamlines and detector setups Deepthi et al. (2015); Fukasawa et al. (2017); Masud et al. (2016); de Gouvêa and Kelly (2016); Deepthi et al. (2018); Blennow et al. (2017); Agarwalla et al. (2016); Masud and Mehta (2016); Coloma and Schwetz (2016); Liao et al. (2016a); Coloma (2016); Denton et al. (2023).
@@
Recent developments in quantum information theory have opened a novel and exciting perspective in the neutrino system Alok et al. (2016); Blasone et al. (2009, 2008); Banerjee et al. (2015); Formaggio et al. (2016); Fu and Chen (2017); Naikoo et al. (2019, 2020); Shafaq and Mehta (2021); Sarkar and Dixit (2021); Blasone et al. (2023); Chattopadhyay and Dighe (2023); Konwar et al. (2024); Dixit et al. (2024); Bouri et al. (2025); Banerjee et al. (2024); Konwar and Yadav (2025); Konwar (2025); Li et al. (2022); Banerjee et al. (2024); Blasone et al. (2010); Li et al. (2022); Blasone et al. (2021); K. Jha et al. (2021); Konwar and Yadav (2024). Specifically, the concept of quantum entanglement, once considered relevant primarily in systems like atomic and optical systems, is now being investigated in the context of fundamental particles, including neutrinos. Several studies have explored the role of quantum entanglement in neutrino oscillations, revealing the presence of nonclassical correlations in flavor evolution. Blasone et al. Blasone et al. (2009) pioneered this approach by treating different neutrino flavor modes as subsystems, showing that flavor oscillations naturally generate bipartite entanglement. Extending this framework, Jha et al. K. Jha et al. (2021) investigated genuine tripartite entanglement in three-flavor oscillations, highlighting how oscillation dynamics encode multipartite quantum correlations. Li et al. Li et al. (2021) examined the interplay between entanglement and measurement uncertainty, demonstrating that uncertainty relations can serve as indicators of quantum correlations in neutrino systems. Subsequent studies by Li et al. Li et al. (2022) and Banerjee et al. Banerjee et al. (2015) further quantified genuine tripartite entanglement and applied quantum information theoretic tools to elucidate the connection between neutrino oscillations. More recently, Konwar and Yadav Konwar and Yadav (2024) analyzed the impact of real NSI for all NSI parameters at a time, on tripartite entanglement, showing that NSI parameters can significantly modify quantum correlations, providing new insights into physics beyond the SM. Together, these works establish neutrino oscillations as a promising platform to study quantum correlations and their effect on both standard and non-standard effects. Entanglement represents a core feature of quantum mechanics, characterising non-classical correlations between different components of a system Horodecki et al. (2009). In neutrino physics, the coherent superposition of mass and flavor eigenstates naturally leads to a form of single-particle mode entanglement Blasone et al. (2008, 2009, 2014), where the different flavor modes of a single neutrino become quantum-mechanically correlated. The key insight is that these entanglement measures can be quantified using neutrino oscillation probabilities, which are directly measurable in neutrino oscillation experiments. To quantify entanglement in neutrino systems, several well-established measures from quantum information theory are employed. In this work, we focus on three such measures: Entanglement of Formation (EOF) Bennett et al. (1996); Guo and Zhang (2020), Concurrence Hill and Wootters (1997); Wootters (1998); Guo and Gour (2019) and Negativity Peres (1996); Vidal and Werner (2002); Sabín and García-Alcaine (2008).
@@
Among the various entanglement measures, the EOF plays a prominent role and captures the minimum average entanglement over all pure state decompositions of a mixed state. Concurrence, initially formulated for bipartite qubit systems, has been successfully extended to tripartite scenarios and provides a robust way to evaluate the degree of entanglement through the reduced density matrices. Another widely used measure is Negativity, which is derived from the Peres-Horodecki criterion Peres (1996). It quantifies entanglement based on the partial transpose of the density matrix, acquiring negative eigenvalues, which serves as a clear signature of non-separability and, consequently, the presence of entanglement in the system. All three measures can be effectively expressed in terms of flavor transition probabilities, providing a framework to study quantum correlations in neutrino oscillations. In this study, we explore the behavior of three key entanglement measures within the framework of three-flavor neutrino oscillations by formulating them in terms of oscillation probabilities. We further examine for first time, how these measures are influenced by the presence of three complex off-diagonal NSI scenarios.
@@
The paper is organized as follows. Section II presents the theoretical formalism underlying neutrino oscillation with NSI, along with a description of mode entanglement in the three-flavor neutrino system. In Section III, we define the three entanglement measures used in our analysis: EOF, Concurrence and Negativity, and express them in terms of neutrino oscillation probabilities. Section IV discusses the experimental description of the DUNE experiment. Section V emphasis the major results coming out from our work. Finally, Section VI summarizes the conclusions of the study.
@@ II Formalism
[…]
and, vice versa, we can also write
[…]
The mixing matrix U, known as the Pontecorvo–Maki–Nakagawa–Sakata (PMNS) matrix, can be parametrized in terms of three mixing angles $\theta_{12}$, $\theta_{23}$ and $\theta_{13}$ and a CP-violating phase $\delta$. The elements of the PMNS matrix are given by
$\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-\iota\delta_{CP}}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}e^{\iota\delta_{CP}}&c_{12}c_{23}-s_{12}s_{13}s_{23}e^{\iota\delta_{CP}}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}e^{\iota\delta_{CP}}&-c_{12}s_{23}-s_{12}s_{13}c_{23}e^{\iota\delta_{CP}}&c_{13}c_{23}\end{pmatrix}$. (4)
Here $s_{ij}=\sin\theta_{ij}$ and $c_{ij}=\cos\theta_{ij}$ with i, j= 1, 2, 3. The time evolution of the flavor state can be written as
[…]
The transition probability for a neutrino of flavor $\alpha$ to be detected as a neutrino of flavor $\beta$ at time $t$ is given by Giunti and Kim (2007):
$P_{(\nu_{\alpha}\rightarrow\nu_{\beta})}(t)=\left|\left<\nu_{\beta}|\nu_{\alpha}(t)\right>\right|^{2}=\left|U^{f}_{\alpha\beta}(t)\right|^{2}$ (7)
In the scenarios beyond the SM, neutrino propagation can be affected by NSI with matter, which introduce additional effective potentials modifying the flavor evolution Hamiltonian. The NSI formalism extends the standard oscillation framework by incorporating these new interactions as perturbations in the matter potential, typically parameterized by dimensionless coefficients $\epsilon_{\alpha\beta}$ with $\alpha,\beta=e,\mu,\tau$. This approach allows for a systematic study of how BSM effects can influence neutrino oscillation probabilities. The widely studied dimension six four-fermion operators responsible for NSIs can be written as Ohlsson (2013); Miranda and Nunokawa (2015); Liao et al. (2017)
$\mathcal{L}_{NSI}^{NC}=2\sqrt{2}G_{F}\sum\limits_{\alpha,\beta,C}\epsilon_{\alpha\beta}^{f,C}(\bar{\nu}_{\alpha}\gamma^{\mu}P_{L}\nu_{\beta})(\bar{f}\gamma_{\mu}P_{C}f),$ (8)
where $G_{F}$ is the Fermi constant, $\epsilon_{\alpha\beta}^{f,C}$ is the parameter which describes the strength of the NSI, $f$ is a first generation SM fermion ($e,u,$ or $d$)), $P_{C}$ denotes the chiral projector with $C=L,R$ corresponding to the left and right-handed chiral projections defined as $P_{L}=(1-\gamma^{5})/2$, and $P_{R}=(1+\gamma^{5})/2$. The indices $\alpha$ and $\beta$ denote the neutrino flavors with $e,\mu,\text{and}~\tau.$
The total Hamiltonian in the presence of matter with NSI can be written as
$\mathcal{H}_{tot}$ $=$ $\mathcal{H}_{vac}+\mathcal{H}_{mat}+\mathcal{H}_{NSI}$ (9)
$\mathcal{H}_{tot}$ $=$ $\begin{pmatrix}E_{1}&0&0\\ 0&E_{2}&0\\ 0&0&E_{3}\end{pmatrix}+U^{\dagger}A\begin{pmatrix}1+\epsilon_{ee}&\epsilon_{e\mu}&\epsilon_{e\tau}\\ \epsilon_{\mu e}&\epsilon_{\mu\mu}&\epsilon_{\mu\tau}\\ \epsilon_{\tau e}&\epsilon_{\tau\mu}&\epsilon_{\tau\tau}\end{pmatrix}U.$ (10)
Where A is the matter potential arising from the coherent forward scattering of electron neutrinos with electrons in matter, and is given by $A=\pm\sqrt{2}G_{F}N_{e}$ with $G_{F}$ is the Fermi constant and $N_{e}$ is the electron number density. The sign of A is positive for neutrinos and negative for antineutrinos. In our analysis, we consider the matter potential as AA= $1.01\times 10^{-13}$ eV, which is associated with Earth’s matter density of $\rho=2.8$ gm/cc. And the NSI parameters $\epsilon_{\alpha\beta}$ are given by
$\epsilon_{\alpha\beta}=\sum_{C,f}\frac{N_{f}}{N_{e}}\epsilon_{\alpha\beta}^{f,C},$ (11)
$\epsilon_{\alpha\beta}^{f,C}$ are the NSI parameters, which describe the interaction strengths between neutrinos of flavors $\alpha$ and $\beta$ with fermions of type $f$, and $N_{f}$ is the number density of $f$ type fermions.
[…]
Here, each term in this expression represents the occupation of one specific flavor mode (e, $\mu$, or $\tau$) while the others remain unoccupied. To characterize entanglement of neutrino states, one constructs the density matrix of the system and evaluates the reduced density matrices after tracing over one or more modes. If the reduced states are mixed, the original state is entangled. The degree of entanglement can be quantified using entanglement measures; EOF, Concurrence, Negativity, etc. In principle, both flavor and mass bases can be used for this characterization, but in this work, we employ the flavor basis. To analyze this entanglement quantitatively, the density matrix of the system at time $t$ is constructed as, $|\rho_{ABC}^{\alpha}(t)=\ket{\nu_{\alpha}(t)}\bra{\nu_{\alpha}(t)}$ where $A$, $B$, and $C$ are the three subsystems of a tripartite system. For the neutrino, these subsystems are chosen as the three flavor modes ($\nu_{e},\nu_{\mu},\nu_{\tau}$) and the density matrix takes the form Banerjee et al. (2024):
$\rho_{e\mu\tau}^{\alpha}(t)=\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&\rho_{22}^{\alpha}&\rho_{23}^{\alpha}&0&\rho_{25}^{\alpha}&0&0&0\\ 0&\rho_{32}^{\alpha}&\rho_{33}^{\alpha}&0&\rho_{35}^{\alpha}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&\rho_{52}^{\alpha}&\rho_{53}^{\alpha}&0&\rho_{55}^{\alpha}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{pmatrix}$, (13)
where the elements of this matrix in the presence of matter and NSI can be represented as.
$\rho_{22}^{\alpha}=\left|\bar{U}^{f}_{\alpha\tau}(t)\right|^{2};~~~~~~~~\rho_{23}^{\alpha}=\bar{U}^{f}_{\alpha\tau}(t)\bar{U}_{\alpha\mu}^{f\ast}(t);~~~~\rho_{25}^{\alpha}=\bar{U}^{f}_{\alpha\tau}(t)\bar{U}_{\alpha e}^{f\ast}(t);$
$\rho_{32}^{\alpha}=\bar{U}^{f}_{\alpha\mu}(t)\bar{U}_{\alpha\tau}^{f\ast}(t);~~~\rho_{33}^{\alpha}=\left|\bar{U}^{f}_{\alpha\mu}(t)\right|^{2};~~~~~~~~~\rho_{35}^{\alpha}=\bar{U}_{\alpha\mu}(t)\bar{U}_{\alpha e}^{f\ast}(t);$
$\rho_{52}^{\alpha}=\bar{U}^{f}_{\alpha e}(t)\bar{U}_{\alpha\tau}^{f\ast}(t);~~~~\rho_{53}^{\alpha}=\bar{U}^{f}_{\alpha e}(t)\bar{U}_{\alpha\mu}^{f\ast}(t);~~~~\rho_{55}^{\alpha}=\left|\bar{U}^{f}_{\alpha e}(t)\right|^{2}.$ (14)
@@
Here $\bar{U}^{f}_{\alpha\beta}(t)=Ue^{-\iota H_{tot}t}U^{-1}$ is the time-evolution operator in the presence of matter and NSI, where $H_{tot}$ is defined in Eq. (10). The probabilities are as follows: $P_{\alpha e}(t)=\left|\bar{U}^{f}_{\alpha e}(t)\right|^{2}$, $P_{\alpha\mu}(t)=\left|\bar{U}^{f}_{\alpha\mu}(t)\right|^{2}$, and $P_{\alpha\tau}(t)=\left|\bar{U}^{f}_{\alpha\tau}(t)\right|^{2}$. In the next section, we will discuss the mathematical expressions of EOF, Concurrence and Negativity in three-flavor neutrino system.